# Emission of Nambu-Goldstone bosons from the semilocal string network

### v1

@@ I Introduction
The cosmic string is a line-like topological defect associated with the breakdown of continuous U(1) symmetry Kibble (1976); Vilenkin and Shellard (2000). The early time evolution of the Universe allows the spontaneous breaking of such a symmetry and then cosmic string forms a network. The network evolves following the so-called scaling regime Kibble (1985); Bennett (1986a, b); Bennett and Bouchet (1990). A remarkable aspect of the cosmic sting is the emission of gravitational waves Vilenkin (1981); Vachaspati and Vilenkin (1985). In particular, the cosmic string associated with the U(1) gauge symmetry breaking predicts the nearly scale-invariant spectrum of gravitational waves, and it is one of the main targets for gravitational wave observations. See Cui et al. (2018); Caprini and Figueroa (2018); Auclair et al. (2020) for reviews.
@@
Furthermore, unlike the Abelian-Higgs model, this model contains massless degrees of freedom associated with Nambu-Goldstone (NG) bosons after the symmetry breaking. Thus, the semilocal string has a channel to decay into those light particles. This is similar to the emission of the (QCD) axions from global (axion) strings, which is extensively studied in the literature Davis (1986); Vilenkin and Vachaspati (1987); Garfinkle and Vachaspati (1987); Yamaguchi et al. (1999); Hagmann et al. (2001); Hiramatsu et al. (2011); Fleury and Moore (2016); Kawasaki et al. (2018); Gorghetto et al. (2018, 2021a); Buschmann et al. (2022); Saikawa et al. (2024); Kim and Son (2025); Benabou et al. (2025). In this scenario, the axion can account for the dark matter as it acquires the mass in the subsequent cosmological evolutions through the non-perturvative effect. Since the topological defects emit particles with low momenta, which is typically the Hubble scale at the emission, the produced particles become non-relativistic and thus behave as cold dark matter. This non-thermal production allows for the dark matter to be very light in contrast with the thermal production.
Similarly, the emission of light gauge bosons is also kinematically allowed in the case with the Abelian-Higgs string if the gauge coupling constant is much smaller than the scalar self-coupling constant which corresponds to the type-II string Long and Wang (2019); Kitajima and Nakayama (2023a, b). In this case, the dark photon can account for the light dark matter.
The light (massless) particle emission is essential in the context of the gravitational wave emissivity from cosmic strings. Notably, if the system allows such light particle emissions, the decay channel to the gravitational wave is significantly suppressed, which drastically changes the resultant gravitational wave spectrum Chang and Cui (2020, 2022); Gorghetto et al. (2021b); Kitajima and Nakayama (2023a). Thus, it is important to investigate the efficiency of the light (massless) particle emission from the string network quantitatively.
@@ II Model
Assuming the Friedmann-Lemetre-Robertson-Waker (FLRW) Universe with zero spatial curvature, one can derive the following system of evolution equations and the constraint equation for the scalar and gauge fields,
@@ III Massless NG mode
Clearly, $\varphi_{r}=\sqrt{\varphi_{1}^{2}+\varphi_{2}^{2}}$ is the radial mode as it feels the potential. The remaining angular component, $\vartheta$, is the direction perpendicular to the radial mode, corresponding to the degree of freedom on the vacuum manifold and thus the massless mode.
The original canonical kinetic terms of the scalar sector can be expressed as [^1] Practically, we mask the region around the string core by multiplying each time-derivative term by $(\varphi_{r}/v)$ as follows, $\dot{\varphi}_{r,{\rm mask}}=\frac{\varphi_{r}}{v}\dot{\varphi}_{r},~~v\dot{\vartheta}_{\rm mask}=\varphi_{r}\dot{\vartheta},~~v\dot{\theta}_{{\pm},{\rm mask}}=\varphi_{r}\dot{\theta}_{\pm}.$ (10)
@@
$=\frac{1}{2}\dot{\varphi}_{r}^{2}+\frac{1}{2}v^{2}(\dot{\theta}_{+}^{2}+\dot{\theta}_{-}^{2}+\dot{\vartheta}^{2})+F(\varphi_{r},\vartheta,\dot{\theta}_{+},\dot{\theta}_{-}),$ (12)
where $F$ is the residual term corresponding to the interaction between 4 modes. Massless modes can be produced via the decay of massive mode through this interaction term.
@@ IV Numerical results
We have performed numerical lattice simulations to solve the nonlinear system of equations (4) and (5) as in previous studies Achucarro et al. (2007); Achúcarro et al. (2014); Lopez-Eiguren et al. (2017). The lattice gauge formulation is adopted in which the discrete gauge symmetry is imposed on each lattice site Moriarty et al. (1988), and the Gauss’s constraint (6) is kept satisfied if it is initially satisfied.
@@
As the setup of our simulations, the number of grid points and the initial boxsize are, respectively, $N^{3}=4096^{3}$ and $L=0.5Nv^{-1}$ ($96v^{-1}$) for the fat (physical) string regime. The model parameters are set as $\lambda=0.025$ and $q=1$, which reads $\beta=0.05$. The conformal time, $\tau$ defined by $dt=ad\tau$, is adopted as the time variable, and we update the field variables using the second-order leap-frog method in our simulations. The stepsize is $0.2\delta x$ with $\delta x=L/N$ being the lattice spacing. The initial conformal time is $v\tau_{i}=1$ and the scale factor is normalized by $a(\tau_{i})=1$. Note that the scale factor is given by $a=v\tau$ in our setup (radiation-dominated universe).
@@
Identifying the semilocal string in the simulation is not trivial because one cannot use the winding number as an indicator of the string core, in contrast to the case with the Abelian-Higgs string. To find the semilocal string, we use the magnetic field concentration in the string Achucarro et al. (2007); Achúcarro et al. (2014). To be more specific, we also perform the simulation with the Abelian-Higgs model to calculate the magnetic field strength averaged over all string segments, $B_{\rm core}^{({\rm AH})}$, whose positions can be identified by the local winding number. Then, the semilocal string can be detected at which the magnetic field strength exceeds the threshold value, $\gamma B_{\rm core}^{({\rm AH})}$ with $\gamma<1$. We set $\gamma=0.5$ in our numerical analysis. Fig. 1 shows snapshots of the evolution of the semilocal string network in our numerical simulations. One can see that strings have endpoints that correspond to global monopolesHindmarsh (1993).
Figure 1: Snapshots of the semilocal string network for the fat string regime. Time evolves from left to right.
To see the scaling behavior of the semilocal string network, we calculate the total length of the string, $\ell_{\rm str}$ with the help of the Abelian-Higgs simulation as follows. The total magnetic field energy, $E_{B}$, can be obtained in both the semilocal and the Abelian-Higgs string cases. Assuming that the magnetic field is confined in the string core, it is given by the product $E_{B}^{\rm(X)}=\mu_{B}^{\rm(X)}\ell_{\rm str}^{\rm(X)}$, where $\mu_{B}$ is the magnetic energy per unit length and X is either SL (semilocal) or AH (Abelian-Higgs). We further assume that the magnetic energy per unit length is equal in both the semilocal and Abelian-Higgs cases, $\mu_{B}^{\rm(SL)}=\mu_{B}^{\rm(AH)}$. Then we obtain $\ell_{\rm str}^{\rm(SL)}=(E_{B}^{\rm(SL)}/E_{B}^{\rm(AH)})\ell_{\rm str}^{\rm(AH)}$. Note that the length of the Abelian-Higgs string can be calculated by using the local winding number.
@@
$\xi=\sqrt{\frac{V}{\ell_{\rm str}}}$ (13)
where $V(=a^{3}L^{3})$ is the physical volume of the simulation box. It increases linearly with time if the network is in the scaling regime. Fig. 2 shows the time evolution of the mean string separation of semilocal strings (red) and Abelian-Higgs strings (blue). The figure clearly demonstrates the scaling behavior of the semilocal string, confirming the results in the literature Achucarro et al. (2007); Achúcarro et al. (2014); Lopez-Eiguren et al. (2017).
Fig. 3 shows the time evolution of the comoving number density of massless NG bosons, $a^{3}n_{\rm NG}$, in the fat/physical string regime (top/bottom panel). The figure demonstrates that the NG bosons are efficiently emitted from the network of semilocal strings. Its linear dependence on the conformal time supports the analytic estimation for the evolution of the number density of emitted NG bosons, which is shown in the next section.
@@
Fig. 4 exhibits the evolution of the spectrum of the comoving number density of the NG mode (the sum of the $\vartheta$ and $\theta_{-}$ modes). It clearly shows that the low-energy peak grows and shifts to lower $k$ as time goes on. Figs. 5 show the spectra for individual massless NG modes, $\vartheta$, $\theta_{-}$, and the radial mode, $\varphi_{r}$, at the final time of the simulation. One can find that low-energy particles are dominantly produced, as in the case with the axion emission from global strings Saurabh et al. (2020); Baeza-Ballesteros et al. (2024) and the light dark photon emission from near-global Abelian-Higgs strings Kitajima and Nakayama (2023a). The peak wave number corresponds to the length scale $\alpha\ell_{H}$ with $\alpha$ and $\ell_{H}$ being, respectively, a numerical constant and the Hubble horizon scale. In both the fat and physical string regimes, we found $\alpha\sim 0.1$. Our results show that the production of the $\theta_{-}$-mode is more efficient than that of the $\vartheta$-mode. It may be due to an artifact from the initial condition, and we expect that two contributions may be equal if we set the initial condition more appropriately.
Figure 4: Evolution of the spectrum of NG modes in the fat string case for $v\tau=29$-10451045 with the interval $v\Delta\tau=64$. Time evolves from bottom to top and the thick black line corresponds to the final time.
Figure 5: Spectra of the comoving number density for two NG modes, $\vartheta$ (solid red) and $\theta_{-}$ (solid blue), and the radial mode, $\varphi_{r}$ (dashed magenta), at the final time of the simulation in the fat/physical string regime (top/bottom panel). The solid curve and shaded region represent respectively the mean value and 1-$\sigma$ error with 10 independent simulations.
V Psendo-NG boson as dark matter
@@
Fig.6 shows the viable parameter region when the pNG boson accounts for the dominant or subdominant component of the dark matter. Note that we have imposed $v<10^{14}$ GeV from the observation of gravitational waves, taking into account the suppression of the gravitational wave emissions due to light particle emissions Kitajima and Nakayama (2023a). This is a rough estimation and detailed study for the gravitational wave emission from the semilocal string network is required, which is left for future work. From Fig.6, the pNG dark matter should be very light, with a lower bound $m_{\rm NG}\gtrsim 10^{-10}$, if the pNG boson is the dominant dark matter. It should be emphasized that the scenario with the formation of the semilocal string network provides an alternative production mechanism of the light pNG dark matter.
@@ VI Discussion
We have numerically shown that massless NG modes are efficiently produced from the semilocal string network. First, we have confirmed the scaling behavior of the network. Then, we have revealed that the comoving number density of emitted NG bosons grows proportional to the Hubble parameter, which is consistent with the analytic estimation. The spectrum of the comoving number density of NG bosons has a peak corresponding to the length scale of $0.1\ell_{H}$ (horizon scale), which implies that a typical momentum of the emitted NG bosons is much smaller than the symmetry breaking scale. If NG bosons acquire the mass due to soft-breaking terms, they can contribute to the dark matter component in the current universe. The mass of such a pNG dark matter can be as small as $10^{-10}$ eV.
@@
In Fig.5, we can see a small peak in the high energy region. Although this feature may be an artifact of the initial condition, it might instead result from the emission at the string endpoints. In those regions, it is possible that the global monopole is accelerated by the string tension, resulting in the emission of relatively high-energy particles. A more quantitative confirmation by numerical simulations is required.

### v2

@@ I Introduction
The cosmic string is a line-like topological defect associated with the breakdown of continuous U(1) symmetry Kibble (1976); Vilenkin and Shellard (2000). The early time evolution of the Universe allows the spontaneous breaking of such a symmetry and then cosmic string forms a network. The network evolves following the so-called scaling regime Kibble (1985); Bennett (1986b, a); Bennett and Bouchet (1990). A remarkable aspect of the cosmic sting is the emission of gravitational waves Vilenkin (1981); Vachaspati and Vilenkin (1985). In particular, the cosmic string associated with the U(1) gauge symmetry breaking predicts the nearly scale-invariant spectrum of gravitational waves, and it is one of the main targets for gravitational wave observations. See Cui et al. (2018); Caprini and Figueroa (2018); Auclair and others (2020) for reviews.
@@
Furthermore, unlike the Abelian-Higgs model, this model contains massless degrees of freedom associated with Nambu-Goldstone (NG) bosons after the symmetry breaking. Thus, the semilocal string has a channel to decay into those light particles. This is similar to the emission of the (QCD) axions from global (axion) strings, which is extensively studied in the literature Davis (1986); Vilenkin and Vachaspati (1987); Garfinkle and Vachaspati (1987); Yamaguchi et al. (1999); Hagmann et al. (2001); Hiramatsu et al. (2011); Fleury and Moore (2016); Kawasaki et al. (2018); Gorghetto et al. (2018, 2021b); Buschmann et al. (2022); Saikawa et al. (2024); Kim et al. (2024); Kim and Son (2025); Benabou et al. (2025). In this scenario, the axion can account for the dark matter as it acquires the mass in the subsequent cosmological evolutions through the non-perturvative effect. Since the topological defects emit particles with low momenta, which is typically the Hubble scale at the emission, the produced particles become non-relativistic and thus behave as cold dark matter. This non-thermal production allows for the dark matter to be very light in contrast with the thermal production.
Similarly, the emission of light gauge bosons is also kinematically allowed in the case with the Abelian-Higgs string if the gauge coupling constant is much smaller than the scalar self-coupling constant which corresponds to the type-II string Long and Wang (2019); Kitajima and Nakayama (2023a, b). In this case, the dark photon can account for the light dark matter. In addition, the Abelian-Higgs string network can also produce axions through the axion-gauge coupling Kitajima and Uwabo-Niibo (2025).
The light (massless) particle emission is essential in the context of the gravitational wave emissivity from cosmic strings. Notably, if the system allows such light particle emissions, the decay channel to the gravitational wave is significantly suppressed, which drastically changes the resultant gravitational wave spectrum Chang and Cui (2020, 2022); Gorghetto et al. (2021a); Kitajima and Nakayama (2023a). Thus, it is important to investigate the efficiency of the light (massless) particle emission from the string network quantitatively.
@@ II Model
Assuming the Friedmann-Lemaître-Robertson-Walker (FLRW) Universe with zero spatial curvature, one can derive the following system of evolution equations and the constraint equation for the scalar and gauge fields,
@@ III Massless NG mode
[…]
The original canonical kinetic terms of the scalar sector can be expressed as
@@
$=\frac{1}{2}\dot{\varphi}_{r}^{2}+\frac{1}{2}\varphi_{r}^{2}(\dot{\theta}_{+}^{2}+\dot{\theta}_{-}^{2}+\dot{\vartheta}^{2})+F(\varphi_{r},\vartheta,\theta_{+},\theta_{-}),$ (14)
where
$F(\varphi_{r},\vartheta,\theta_{+},\theta_{-})=\varphi_{r}^{2}\cos(2\vartheta)\dot{\theta}_{+}\dot{\theta}_{-}.$ (15)
In practice, we first compute the time derivatives of $\varphi_{c}$ and $\theta_{c}$ as
$\dot{\varphi}_{c}=\frac{2{\rm Re}(\Phi_{c}^{*}\dot{\Phi}_{c})}{\varphi_{c}}\,,~~\dot{\theta}_{c}=\frac{2{\rm Im}(\Phi_{c}^{*}\dot{\Phi}_{c})}{\varphi_{c}^{2}}\,.$ (16)
Then, we obtain $\dot{\varphi}_{r},\dot{\vartheta}$, and $\dot{\theta}_{\pm}$ as
$\dot{\varphi}_{r}=\frac{\varphi_{1}\dot{\varphi}_{1}+\varphi_{2}\dot{\varphi}_{2}}{\varphi_{r}}\,$, (17)
$\dot{\vartheta}=\frac{\varphi_{1}\dot{\varphi}_{2}-\varphi_{2}\dot{\varphi}_{1}}{\varphi_{r}^{2}}\,$, (18)
$\dot{\theta}_{\pm}=\frac{\dot{\theta}_{1}\pm\dot{\theta}_{2}}{2}\,$, (19)
which allows us to compute the energy density and the number density of each mode [^2] Practically, we mask the region around the string core by multiplying each time-derivative term by $(\varphi_{r}/v)$ as follows, $\dot{\varphi}_{r,{\rm mask}}=\frac{\varphi_{r}}{v}\dot{\varphi}_{r},~~v\dot{\vartheta}_{\rm mask}=\varphi_{r}\dot{\vartheta},~~v\dot{\theta}_{{\pm},{\rm mask}}=\varphi_{r}\dot{\theta}_{\pm}.$ (20) Similarly, we also mask the point at which either $\varphi_{1}$ or $\varphi_{2}$ vanishes. .
Since the semilocal string can be regarded as a linear excitation of $\varphi_{r}$, massless modes can be emitted from the strings via the above interaction terms. However, since the dynamics is fully nonlinear, the perturbative analysis cannot be applied in this case  [^3] The NG boson emission from the global string can be described by the effective Kalb-Ramond action. See the discussion in Sec. VI. . Indeed, the massless NG modes can be excited by the fully nonlinear dynamics of semilocal strings extended over the horizon scale. Naively, the oscillation of horizon-sized loop can excite the NG modes with horizon-scale wavelength, but we need precise numerical simulations to follow such a nonlinear emission process.
@@ IV Numerical results
We have performed numerical lattice simulations to solve the nonlinear system of Eqs. (4) and (5) as in previous studies Achucarro et al. (2007); Achúcarro et al. (2014); Lopez-Eiguren et al. (2017). The lattice gauge formulation is adopted in which the discrete gauge symmetry is imposed on each lattice site Moriarty et al. (1988), and the Gauss’s constraint (6) is kept satisfied if it is initially satisfied.
@@
As the setup of our simulations, the number of grid points and the initial boxsize are, respectively, $N^{3}=4096^{3}$ and $L=0.5Nv^{-1}$ ($96v^{-1}$) for the fat (physical) string regime. The model parameters are set as $\lambda=0.025$ and $q=1$, which reads $\beta=0.05$.  [^4] To produce a sufficiently populated network, small values of $\beta$ are favored Achucarro et al. (2007). However, for very small $\beta$, the width of the magnetic flux and the scalar excitation become too different in scale. Therefore, in this work we adopt the range of beta used in Ref. Achúcarro et al. (2014). We show the results for another value of $\beta$ in App. A, which confirm that the qualitative behavior do not depends on $\beta$. The conformal time, $\tau$ defined by $dt=ad\tau$, is adopted as the time variable, and we update the field variables using the second-order leap-frog method in our simulations. The stepsize is $0.2\delta x$ with $\delta x=L/N$ being the lattice spacing. The initial conformal time is $v\tau_{i}=1$ and the scale factor is normalized by $a(\tau_{i})=1$. Note that the scale factor is given by $a=v\tau$ in our setup (radiation-dominated universe).
@@
Identifying the semilocal string in the simulation is not trivial because one cannot use the winding number as an indicator of the string core, in contrast to the case with the Abelian-Higgs string. To find the semilocal string, we use the magnetic field concentration in the string Achucarro et al. (2007); Achúcarro et al. (2014). To be more specific, we also perform the simulation with the Abelian-Higgs model to calculate the magnetic field strength averaged over all string segments, $B_{\rm core}^{({\rm AH})}$, whose positions can be identified by the local winding number. Then, the semilocal string can be detected at which the magnetic field strength exceeds the threshold value, $\gamma B_{\rm core}^{({\rm AH})}$ with $\gamma<1$. We set $\gamma=0.5$ in our numerical analysis. Fig. 1 shows snapshots of the evolution of the semilocal string network in our numerical simulations. One can see that strings have endpoints that correspond to global monopolesHindmarsh (1993). Interaction between monopoles leads to the merging of segments, whereas small segments shrink and disappear due to the string tension. By the end of the simulation, almost all segments have vanished from the simulation box except for those larger than the horizon scale.
Figure 1: Snapshots of the semilocal string network for the fat string regime. Time evolves from top left to top right and bottom left to bottom right, corresponding to the conformal time $\tau=61v^{-1}$-$261v^{-1}$ with the interval $\Delta\tau=40v^{-1}$.
To see the scaling behavior of the semilocal string network, we calculate the total length of the string, $\ell_{\rm str}$ with the help of the Abelian-Higgs simulation as follows. The total magnetic field energy, $E_{B}$, can be obtained in both the semilocal and the Abelian-Higgs string cases. Assuming that the magnetic field is confined in the string core, it is given by the product $E_{B}^{\rm(X)}=\mu_{B}^{\rm(X)}\ell_{\rm str}^{\rm(X)}$, where $\mu_{B}$ is the magnetic energy per unit length and X is either SL (semilocal) or AH (Abelian-Higgs). For a straight static string, $g(r)$ in Eq. (3) is identical to that of the Abelian-Higgs string when $\beta$ is the same Vachaspati and Achucarro (1991). Then we take $\mu_{B}^{\rm(SL)}=\mu_{B}^{\rm(AH)}$ and obtain $\ell_{\rm str}^{\rm(SL)}=(E_{B}^{\rm(SL)}/E_{B}^{\rm(AH)})\ell_{\rm str}^{\rm(AH)}$. Note that the length of the Abelian-Higgs string can be calculated by using the local winding number.
@@
$\xi_{\rm phys}=\sqrt{\frac{V}{\ell_{\rm str}}}$ (21)
where $V(=a^{3}L^{3})$ is the physical volume of the simulation box. It increases linearly with time if the network is in the scaling regime. Fig. 2 shows the time evolution of the comoving mean string separation, $\xi=\xi_{\rm phys}/a$, of semilocal strings (red) and Abelian-Higgs strings (blue). The figure clearly demonstrates the scaling behavior of the semilocal string, confirming the results in the literature Achucarro et al. (2007); Achúcarro et al. (2014); Lopez-Eiguren et al. (2017). The oscillatory behavior of $\xi$ at early $\tau$ in the semilocal string is a consequence of spurious transient oscillations of the background magnetic field induced by the initial configuration of $\phi$ since the simulation starts with the zero gauge fields. As $\tau$ increases, these artificial oscillations are damped, and we can observe the scaling behavior of the network at late times.
Masless NG bosons can be emitted from the network of semilocal strings, mainly through the decay of loops. Initially, the loop has a horizon size, and its oscillations can excite the NG modes with Hubble-scale frequencies. Fig. 3 shows the time evolution of the comoving number density of massless NG bosons, $a^{3}n_{\rm NG}$, in the fat/physical string regime (top/bottom panel). The figure demonstrates that the NG bosons are efficiently emitted from the network of semilocal strings. Its linear dependence on the conformal time supports the analytic estimation with the assumption that the emitted NG bosons has Hubble-scale frequencies, which is shown in the next section.
@@
Fig. 4 exhibits the evolution of the spectrum of the comoving number density of the NG mode (the sum of the $\vartheta$ and $\theta_{-}$ modes). It clearly shows that the low-energy peak grows and shifts to lower comoving momentum $k$ as time goes on. In particular, the peak physical momentum is proportional to the Hubble parameter ($\bar{k}_{\rm peak}\sim 10H$) when the network evolves in the scaling regime (see the dashed vertical lines). Note that the gradual shift of the peak wavenumber implies that the emission is not due to the decay of heavy modes. We can also see a small peak in the high-$k$ region, which may be an artifact caused by the finite resolution.
Figs. 5 show the spectra for individual massless NG modes, $\vartheta$, $\theta_{-}$, and the radial mode, $\varphi_{r}$, at the final time of the simulation. One can find that low-energy particles are dominantly produced, as in the case with the axion emission from global strings Saurabh et al. (2020); Baeza-Ballesteros et al. (2024) and the light dark photon emission from near-global Abelian-Higgs strings Kitajima and Nakayama (2023a). The peak wave number corresponds to the length scale $\alpha\ell_{H}$ with $\alpha$ and $\ell_{H}$ being, respectively, a numerical constant and the Hubble horizon scale. In both the fat and physical string regimes, we found $\alpha\sim 0.1$. Our results show that the production of the $\theta_{-}$-mode is more efficient than that of the $\vartheta$-mode. We attribute this difference to the asymmetric couplings of the massless NG modes to $\varphi_{r}$ as in Eq. (15). However, this asymmetry does not affect the qualitative nature of our conclusion.
Figure 4: Evolution of the spectrum of NG modes in the fat string case for $v\tau=85,\,149,\,277,\,533,\,1045$. The $k$ on the horizontal axis denotes the comoving momentum. Time evolves from bottom to top and the thick black line corresponds to the final time. The dashed magenta and black lines correspond to $\bar{k}=10H$ at $v\tau=533,\,1045$, respectively.
Figure 5: Spectra of the comoving number density for two NG modes, $\vartheta$ (solid red) and $\theta_{-}$ (solid blue), and the radial mode, $\varphi_{r}$ (dashed magenta), at the final time of the simulation in the fat/physical string regime (top/bottom panel). The solid curve and shaded region represent respectively the mean value and 1-$\sigma$ error with 10 independent simulations. The dashed line corresponds to $\bar{k}=10H$ at that time.
V Pseudo-NG boson as dark matter
@@
Fig.6 shows the viable parameter region when the pNG boson accounts for the dominant or subdominant component of the dark matter. Note that we have imposed $v<10^{14}\,{\rm GeV}$ from the observation of gravitational waves, taking into account the suppression of the gravitational wave emissions due to light particle emissions Kitajima and Nakayama (2023a). This is a rough estimation and detailed study for the gravitational wave emission from the semilocal string network is required, which is left for future work. From Fig.6, the pNG dark matter should be very light, with a lower bound $m_{\rm NG}\gtrsim 10^{-10}\,{\rm eV}$, if the pNG boson is the dominant dark matter. It should be emphasized that the scenario with the formation of the semilocal string network provides an alternative production mechanism of the light pNG dark matter.
@@ VI Discussion
We have numerically shown that massless NG modes are efficiently produced from the semilocal string network. First, we have confirmed the scaling behavior of the network. Then, we have revealed that the comoving number density of emitted NG bosons grows proportional to the Hubble parameter, which is consistent with the analytic estimation. The spectrum of the comoving number density of NG bosons has a peak corresponding to the length scale of $0.1\ell_{H}$ (horizon scale), which implies that a typical momentum of the emitted NG bosons is much smaller than the symmetry breaking scale. If NG bosons acquire the mass due to soft-breaking terms, they can contribute to the dark matter component in the current universe. The mass of such a pNG dark matter can be as small as $10^{-10}\,{\rm eV}$.
@@
Our simulation results show that the characteristic momentum of the emitted NG bosons is of order the Hubble scale. This behavior is analogous to the axion radiation from a global string, which is successfully described by a Kalb–Ramond field coupled to a Nambu–Goto string Kalb and Ramond (1974); Vilenkin and Vachaspati (1987). Motivated by this analogy, we expect that the NG-boson emission from a semilocal string can likewise be captured by antisymmetric tensor fields dual to $\theta_{-}$ and $\vartheta$. However, it is not obvious that these tensor fields couple to the string in the same manner as in the Kalb–Ramond action, since $\theta_{-}$ and $\vartheta$ do not wind around the string core. We also note that the doublet $n$ in Eq. (3) can undergo dynamical fluctuations near the string core, which may provide an alternative mechanism for coupling the NG modes to the string. Clarifying the structure of the corresponding low-energy theory remains an open question.